OpenStudy (anonymous):

Express the complex number in trigonometric form. -2 + 2square root of threei

OpenStudy (anonymous):

@mathmath333

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

SOMEONE HELP

OpenStudy (amistre64):

what 2 things do we need to determine for the trig form?

OpenStudy (anonymous):

the absolute value? @amistre64

OpenStudy (amistre64):

hmm, i was thinking a radius and an angle

OpenStudy (anonymous):

oh sorry, I have no idea how to do this

OpenStudy (amistre64):

spose we plot some point (x,y) define as some complex number: x + iy  we can then determine the distance from the origin, and the angle the ray makes with the x axis

OpenStudy (amistre64):

r = sqrt(x^2+y^2) a = tan^(-1) (y/x)

OpenStudy (anonymous):

okay so I plug the numbers into that formula?

OpenStudy (amistre64):

what else would you do with a formula? trig defines x = r cos(a) and y=r sin(a) therefore our trig form of a complex number is: x + iy to r cos(a) + r i sin(a) or simplified to: r (cos(a) + i sin(a))

OpenStudy (anonymous):

4(cos5pi/6+isin5pi/6)?

OpenStudy (anonymous):

?

OpenStudy (amistre64):

\[\Large \underbrace{-2}_x+i~\underbrace{2\sqrt3}_y\] \[r=\sqrt{x^2+y^2}=\sqrt{4+12}=4~;~is~fine\] \[\alpha=tan^{-1}(-\frac{1}{\sqrt3})=30^o~(Q2)\]

OpenStudy (amistre64):

yeah, looks good to me in my head lol

OpenStudy (amistre64):